please help me solve this equation4(1-3x)+7x=9

Answer by checkley71(8403)

4(1-3x)+7x=9 or 4-12x+7x=9 or -5x=5 or x=-1
proof: 4(1-3(-1))+7(-1)=9 or 4(1+3)-7=9 or 4(4)-7=9 or 16-7=9 or 9=9.